# MixFlow: Mixture-Conditioned Flow Matching for Out-of-Distribution Generalization

### v1

@@ Abstract
Achieving robust generalization under distribution shift remains a central challenge in conditional generative modeling, as existing conditional flow-based methods often struggle to extrapolate beyond the training conditions. We introduce MixFlow, a conditional flow-matching framework for descriptor-controlled generation that directly targets this limitation by jointly learning a descriptor-conditioned base distribution and a descriptor-conditioned flow field via shortest-path flow matching. By modeling the base distribution as a learnable, descriptor-dependent mixture, MixFlow enables smooth interpolation and extrapolation to unseen conditions, leading to substantially improved out-of-distribution generalization. We provide analytical insights into the behavior of the proposed framework and empirically demonstrate its effectiveness across multiple domains, including prediction of responses to unseen perturbations in single-cell transcriptomic data and high-content microscopy–based drug screening tasks. Across these diverse settings, MixFlow consistently outperforms standard conditional flow-matching baselines. Overall, MixFlow offers a simple yet powerful approach for achieving robust, generalizable, and controllable generative modeling across heterogeneous domains.
@@ 1 Introduction
This challenge of out-of-distribution (OOD) generalization***In this manuscript the term OOD generalization refers to the ability of a model to take in an unseen descriptor and accurately generate the corresponding population. This departs from the common notion of the term, that is how a distribution shift in a predictor’s input instances affects its predictions. Moreover, in our setting the unseen descriptor should still come from the same distribution. is particularly acute in scientific domains. Consider drug discovery, where one wishes to predict cellular responses to novel compounds (Subramanian et al., 2017; Srivatsan et al., 2020), or genetic screening, where the goal is to anticipate phenotypic effects of unseen perturbations (Norman et al., 2019; Replogle et al., 2022). In such settings, the space of possible conditions (drugs, genetic perturbations, material compositions) is combinatorially vast, and any practical model must extrapolate beyond its training distribution. Recent work has highlighted that even sophisticated deep learning approaches often fail to outperform simple baselines when evaluated on truly held-out conditions (Ahlmann-Eltze et al., 2025), underscoring the difficulty of this problem.
@@
Recent work has begun to explore alternatives to fixed base distributions. Chen et al. (2025) propose modelling the velocity field as a Gaussian mixture to improve few-step sampling, though this does not address the base distribution itself. Atanackovic et al. (2024) introduce Meta Flow Matching, which conditions on initial populations via graph neural network embeddings to model dynamics on the Wasserstein manifold. However, this method works under the assumption that the starting distribution is given to the model and that the nature of the difference between these observed initial distributions is the main source of variability between different effects. In practice, this model is designed to learn how to amortise the effects across contexts rather than generalise to new perturbations. Other approaches design informative priors for specific applications to shorten transport paths and improve sampling efficiency (Bunne et al., 2023). However, none of these methods jointly learn a descriptor-conditioned base distribution with the explicit goal of improving OOD generalization, nor do they provide theoretical justification for why this design choice matters.
Figure 1: Standard conditional flow matching versus MixFlow. (a) Conventional approaches fix the base distribution to a standard Gaussian and condition only the velocity field on the descriptor $y$, effectively learning independent flows for each condition. (b) MixFlow conditions both the base distribution (as a Gaussian mixture) and the velocity field on the descriptor. Similar conditions produce similar base distributions, requiring only small flow corrections and enabling smooth extrapolation to unseen conditions.
@@
We introduce MixFlow, a conditional flow-matching framework that jointly learns a descriptor-conditioned Gaussian mixture base distribution and a descriptor-conditioned velocity field trained via shortest-path (optimal transport) flow matching. The key architectural change is simple: instead of sampling from a fixed Gaussian, MixFlow uses learned networks to predict condition-specific mixture parameters that define the base distribution. Samples from this adaptive base are then transported to the target via a learned velocity field. Because similar descriptors produce similar base distributions, the model can smoothly interpolate and extrapolate to unseen conditions, a capability that fixed-base approaches fundamentally lack.
@@
We evaluate MixFlow across diverse domains to demonstrate its broad applicability. On synthetic benchmarks involving rotated letter silhouettes inspired by Atanackovic et al. (2024), MixFlow consistently outperforms conditional flow matching when generalizing to unseen rotations. On image-based phenotypic assays (Caie et al., 2010; Sypetkowski et al., 2023; Bray et al., 2016), MixFlow substantially reduces distributional distances compared to baselines when predicting cellular morphologies under novel compound treatments. On single-cell transcriptomic datasets (Norman et al., 2019; Lotfollahi et al., 2023; Replogle et al., 2022), MixFlow achieves strong performance in predicting gene expression responses to unseen genetic and chemical perturbations.
@@ 1.1 Contributions
  • We demonstrate consistent improvements over conditional flow matching across eight datasets spanning synthetic, image-based, and transcriptomic domains, establishing MixFlow as a flexible framework for conditional generation with improved OOD generalization.
@@ 1.2 Related Work
Gaussian mixture models (GMMs) have a long history as flexible density estimators with universal approximation properties (Nguyen, 2019; Zheng and Tang, 2020). Recent work has revisited GMMs in the context of modern generative models. Chen et al. (2025) propose GMFlow, which models the velocity distribution as a Gaussian mixture to capture multimodal flow directions, improving few-step sampling and reducing over-saturation artifacts. Ahamed et al. (2025) use GMMs in latent spaces for molecular generation. These approaches modify how the velocity or score is represented but do not condition the base distribution on input descriptors. In contrast, MixFlow learns a descriptor-conditioned GMM as the source distribution, explicitly encoding condition similarity structure to improve OOD generalization.
@@
Several works have explored moving beyond fixed Gaussian priors. In the context of variational autoencoders, learned priors have been used to improve expressiveness (Bereket and Karaletsos, 2023). For flow-based models, Bunne et al. (2023) learn transport maps via neural optimal transport for single-cell perturbation modeling, though with a focus on paired control-treatment settings rather than OOD generalization. Meta Flow Matching (Atanackovic et al., 2024) conditions on initial populations by embedding them via graph neural networks, enabling generalization across different starting distributions on the Wasserstein manifold. This approach is complementary to ours: while Meta Flow Matching learns to adapt to varying initial populations, MixFlow learns to predict appropriate base distributions from condition descriptors, directly targeting the OOD generalization problem.
@@
OOD generalization remains a central challenge in machine learning, with extensive work on domain adaptation, invariant learning, and distributionally robust optimization. In the context of perturbation biology, recent analyses have shown that deep learning models often fail to generalize beyond training perturbations, sometimes underperforming simple linear baselines (Ahlmann-Eltze et al., 2025). This has motivated the development of methods that explicitly encode biological structure, such as pathway information (Roohani et al., 2024) or chemical similarity (Subramanian et al., 2017; Wang et al., 2016). MixFlow takes a complementary approach by encoding similarity structure directly into the generative model architecture through the learned base distribution, providing a general mechanism for OOD generalization that does not require domain-specific inductive biases.
@@ 2.1 Background on Conditional Flow Matching
3 MixFlow: Mixture-Conditioned Flow Matching
Standard conditional flow matching fixes the base distribution to $\mathcal{N}(\mathbf{0}_{D},\sigma^{2}\mathbf{I}_{D\times D})$ and conditions only the decoder, effectively yielding independent estimators for each population and limiting generalization (See, Fig. 1(a)). MixFlow instead conditions both the base distribution and the flow module on perturbations and covariates (Fig. 1(b)).
@@ 3.1 Training objectives
MixFlow is trained with two complementary losses.
@@ 3.2 Inference
At test time, given a novel descriptor $\bm{y}_{test}$, MixFlow predicts the base distribution $\mu_{\mathrm{GMM}(h_{\bm{\Theta}}(\bm{y}_{test}),h_{\bm{p}}(\bm{y}_{test}))}$ and evolves it with $v(.,t;\bm{y}_{test}),\;t\in[0,1]$ to generate samples approximating $\rho_{test}$.
Algorithm 1 Training MixFlow with mixture-conditioned flow matching
@@ 5 Experiments
We evaluate MixFlow against Vanilla Conditional Flow Matching (Vanilla CFM) on a set of synthetic and experimental datasets. Our goal is to assess MixFlow’s generalizability and its robustness across multiple conditions, datasets, and hyperparameters. We seek empirical evidence of the need for the Gaussian Mixture Model (GMM) to describe the base distribution and its impact on the distributional distance between the inferred final distributions and the empirical ones.
5.1 MixFlow on Synthetic Data
@@
For MixFlow, we construct a mixture-conditioned base for each population. We set the number of base modes $I$ equal to the number of conditions and assign one Gaussian component per condition. In our synthetic setup, we do not use $h_{\bm{\Theta}}(\bm{y}_{i})$ to predict the component means; instead, the GMM mode locations are implemented as free learnable parameters, each initialized from a single rendered sample of its corresponding condition. The mixture weights are still predicted by the MLP $h_{\bm{p}}(\bm{y}_{i})$. The velocity field $v(\cdot,t;\bm{y}_{i})$ is then trained using the flow-matching objectives in Eqs. 67 to transport the resulting $\mu_{\mathrm{GMM}(\bm{\Theta}^{(i)},h_{\bm{p}}(\bm{y}_{i}))}$ toward $\rho_{i}$.
@@
As a baseline, CFM fixes the base to $\mathcal{N}(\mathbf{0}_{D},\mathbf{I}_{D\times D})$ and samples $x_{0}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ at generation time. We train both MixFlow and CFM on a synthetic dataset of 6 letters with $R=20$ rotations per letter, and study rotation interpolation generalization. Specifically, we generate $R$ rotation angles uniformly spaced in $0,2\pi]$ and split them by index: the training set contains the 10 even-indexed rotations, while the held-out set contains the 10 odd-indexed rotations. The validation set is formed by selecting a single val_letter and evaluating that letter only on the held-out (odd-indexed) rotations; all letters (including val_letter) are present in training but only at the even-indexed rotations. We report two independent runs with identical settings except for the choice of validation letter, using $\texttt{val\_letter}\in\{\texttt{S},\texttt{H}\}$. We do this to check that conclusions are not tied to a single character’s geometry and to reduce sensitivity to a particular letter choice. In each run, the held-out set contains 10 unseen rotations (odd indices) for the chosen validation letter (S or H). Results are reported in Table 1. Qualitative samples are available in Figure 2.
We evaluate both models using three standard distributional metrics: Wasserstein-1 (W1), Wasserstein-2 (W2), and Maximum Mean Discrepancy (MMD). For each experiment, we select the best checkpoint based on validation Wasserstein-2 distance and report all metrics at this checkpoint. Across both validation-letter settings (S and H), MixFlow consistently improves over conditional flow matching (CFM) in all three metrics. In particular, MixFlow achieves lower W1 (reducing by $\sim 39\%$ and $\sim 13\%$), lower W2 (reducing by $\sim 16\%$ and $\sim 13\%$), and lower MMD (reducing by $\sim 17\%$ and $\sim 12\%$) for both experiments, demonstrating that mixture-conditioned base distributions provide a more expressive and accurate starting point than the fixed isotropic Gaussian used by CFM.
@@
Figure 2: Qualitative comparison of Conditional Flow Matching (CFM) and MixFlow on the synthetic letter–rotation benchmark. Each block shows samples at different integration times ($t=0$, $t=0.5$, $t=1$) and the corresponding target distribution. MixFlow produces smoother trajectories and more faithful reconstructions on both training and unseen rotations, while CFM exhibits drift and shape distortion, particularly on validation conditions.
Dataset Model MMD $\downarrow$ W1 $\downarrow$ W2 $\downarrow$
Letter S CFM 0.689 $\pm$ 0.0140 0.01533 $\pm$ 0.00094 0.5667 $\pm$ 0.0114
MixFlow 0.5737 $\pm$ 0.0071 0.00933 $\pm$ 0.00047 0.4723 $\pm$ 0.0057
Letter H CFM 0.8363 $\pm$ 0.0193 0.1857 $\pm$ 0.0061 0.6777 $\pm$ 0.0168
MixFlow 0.7297 $\pm$ 0.0061 0.1380 $\pm$ 0.0016 0.5867 $\pm$ 0.0024
5.2 Using MixFlow to predict single-cell responses for unseen interventions

### v2

# Shortest-Path Flow Matching with Mixture-Conditioned Bases for OOD Generalization to Unseen Conditions

@@ Abstract
Robust generalization under distribution shift remains a key challenge for conditional generative modeling: conditional flow-based methods often fit the training conditions well but fail to extrapolate to unseen ones. We introduce SP-FM, a shortest-path flow-matching framework that improves out-of-distribution (OOD) generalization by conditioning both the base distribution and the flow field on the condition. Specifically, SP-FM learns a condition-dependent base distribution parameterized as a flexible, learnable mixture, together with a condition-dependent vector field trained via shortest-path flow matching. Conditioning the base allows the model to adapt its starting distribution across conditions, enabling smooth interpolation and more reliable extrapolation beyond the observed training range. We provide theoretical insights into the resulting conditional transport and show how mixture-conditioned bases enhance robustness under shift. Empirically, SP-FM is effective across heterogeneous domains, including predicting responses to unseen perturbations in single-cell transcriptomics and modeling treatment effects in high-content microscopy–based drug screening. Overall, SP-FM provides a simple yet effective plug-in strategy for improving conditional generative modeling and OOD generalization across diverse domains.
@@ 1 Introduction
Out-of-distribution (OOD) generalization is particularly challenging in scientific applications because models are routinely asked to generate responses for conditions that were never observed during training. Here, we use condition to mean the intervention or experimental setting that indexes a data-generating distribution—for example a drug identity, a genetic perturbation, or a specific rotation of a letter—rather than a shift in the input instances of a predictor. Conditional generative models are often brittle in this regime: they may interpolate within the training condition set but degrade sharply on truly held-out conditions, producing overly averaged samples or missing condition-specific effects.***In this paper, OOD generalization refers to generating from an unseen condition (intervention/setting) excluded from training, not to test-time covariate shift in predictor inputs. We assume the unseen condition is drawn from the same overall condition distribution.
This issue appears across scientific domains. For example, in drug discovery one aims to predict cellular responses to novel compounds (Subramanian et al., 2017; Srivatsan et al., 2020). Similarly, in genetic screening the goal is to predict phenotypic effects of unseen perturbations (Norman et al., 2019; Replogle et al., 2022). In such settings, the condition space (e.g., drugs, perturbations, compositions and their combinations) is combinatorially large, making exhaustive coverage impossible and forcing practical models to extrapolate beyond the training conditions. Consistent with this, recent benchmarks report that even sophisticated deep learning approaches often fail to outperform simple baselines when evaluated on truly held-out conditions (Ahlmann-Eltze et al., 2025), underscoring the difficulty of robust conditional generation under shift.
@@
Recent work has begun to move beyond fixed base distributions, but along directions that differ from our objective. Chen et al. (2025) model the velocity field as a Gaussian mixture to enable efficient few-step sampling, without learning a condition-dependent base distribution. Atanackovic et al. (2024) propose Meta Flow Matching, which conditions the dynamics on an observed initial population encoded by a graph neural network, effectively assuming access to the starting distribution at test time and focusing on transferring dynamics across contexts. Other approaches construct application-specific priors to shorten transport paths and improve sampling efficiency (Bunne et al., 2023). In contrast, SP-FM explicitly learns a condition-dependent base distribution jointly with the flow field, with the goal of improving OOD generalization to unseen conditions, and provides theoretical motivation for why conditioning the base can matter.
Figure 1: Standard conditional flow matching versus SP-FM. (a) Conventional approaches fix the base distribution to a standard Gaussian and condition only the velocity field on the descriptor $y$, effectively learning independent flows for each condition. (b) SP-FM conditions both the base distribution (as a Gaussian mixture) and the velocity field on the descriptor. For each target distribution, the base distribution is selected such that the length of the geodesic connecting the two is minimised. Similar conditions produce similar base distributions, requiring only small flow corrections and enabling smooth extrapolation to unseen conditions.
@@
We introduce SP-FM, a conditional flow-matching framework that jointly learns a descriptor-conditioned Gaussian mixture base distribution and a descriptor-conditioned velocity field trained via shortest-path (optimal transport) flow matching. The key architectural change is simple: instead of sampling from a fixed Gaussian, SP-FMuses learned networks to predict condition-specific mixture parameters that define the base distribution. Samples from this adaptive base are then transported to the target via a learned velocity field. Because similar descriptors produce similar base distributions, the model can smoothly interpolate and extrapolate to unseen conditions, a capability that fixed-base approaches fundamentally lack.
@@
While conditional flow matching is often paired with domain-specific architectures and inductive biases (e.g., in single-cellKlein et al. (2025) and morphology-based perturbation predictionNavidi et al. (2024); Zhang et al. (2025b)), our goal is a general improvement that can be plugged into existing pipelines. We therefore compare primarily to standard conditional flow-matching baselines under matched capacity and training budgets to avoid confounding gains from task-specific design choices. On synthetic rotated-letter benchmarks (Atanackovic et al., 2024), SP-FM improves generalization to unseen rotations. On image-based phenotypic assays (Caie et al., 2010; Sypetkowski et al., 2023; Bray et al., 2016), it reduces distributional distances when predicting morphologies under novel compound treatments. On single-cell transcriptomic datasets (Norman et al., 2019; Lotfollahi et al., 2023; Replogle et al., 2022), it accurately predicts gene-expression responses to unseen genetic and chemical perturbations.
@@ 1.1 Contributions
  • We demonstrate consistent improvements over conditional flow matching across eight datasets spanning synthetic, image-based, and transcriptomic domains, establishing SP-FMas a flexible framework for conditional generation with improved OOD generalization.
@@ 1.2 Related Work
Gaussian mixture models (GMMs) have a long history as flexible density estimators with universal approximation properties (Nguyen, 2019; Zheng and Tang, 2020). Recent work has revisited GMMs in the context of modern generative models. Chen et al. (2025) propose GMFlow, which models the velocity distribution as a Gaussian mixture to capture multimodal flow directions, improving few-step sampling and reducing over-saturation artifacts. Ahamed et al. (2025) use GMMs in latent spaces for molecular generation. These approaches modify how the velocity or score is represented but do not condition the base distribution on input descriptors. In contrast, SP-FMlearns a descriptor-conditioned GMM as the source distribution, explicitly encoding condition similarity structure to improve OOD generalization.
@@
Several works have explored moving beyond fixed Gaussian priors. In the context of variational autoencoders, learned priors have been used to improve expressiveness (Bereket and Karaletsos, 2023). For flow-based models, Bunne et al. (2023) learn transport maps via neural optimal transport for single-cell perturbation modeling, though with a focus on paired control-treatment settings rather than OOD generalization. Meta Flow Matching (Atanackovic et al., 2024) conditions on initial populations by embedding them via graph neural networks, enabling generalization across different starting distributions on the Wasserstein manifold. This approach is complementary to ours: while Meta Flow Matching learns to adapt to varying initial populations, SP-FMlearns to predict appropriate base distributions from condition descriptors, directly targeting the OOD generalization problem.
@@
OOD generalization remains a central challenge in machine learning, with extensive work on domain adaptation, invariant learning, and distributionally robust optimization. In the context of perturbation biology, recent analyses have shown that deep learning models often fail to generalize beyond training perturbations, sometimes underperforming simple linear baselines (Ahlmann-Eltze et al., 2025). This has motivated the development of methods that explicitly encode biological structure, such as pathway information (Roohani et al., 2024) or chemical similarity (Subramanian et al., 2017; Wang et al., 2016). SP-FMtakes a complementary approach by encoding similarity structure directly into the generative model architecture through the learned base distribution, providing a general mechanism for OOD generalization that does not require domain-specific inductive biases.
@@ 2.1 Background on Conditional Flow Matching
3 Mixture-Conditioned Flow Matching
Standard conditional flow matching fixes the base distribution to $\mathcal{N}(\mathbf{0}_{D},\sigma^{2}\mathbf{I}_{D\times D})$ and conditions only the decoder, effectively yielding independent estimators for each population and limiting generalization (See, Fig. 1(a)). SP-FM instead conditions both the base distribution and the flow module on perturbations and covariates (Fig. 1(b)).
@@ 3.1 Training objectives
SP-FM is trained with two complementary losses.
@@ 3.2 Inference
At test time, given a novel descriptor $\bm{y}_{test}$, SP-FM predicts the base distribution $\mu_{\mathrm{GMM}(h_{\bm{\Theta}}(\bm{y}_{test}),h_{\bm{p}}(\bm{y}_{test}))}$ and evolves it with $v(.,t;\bm{y}_{test}),\;t\in[0,1]$ to generate samples approximating $\rho_{test}$.
Algorithm 1 Training SP-FM with mixture-conditioned flow matching
@@ 5 Experiments
We evaluate SP-FM against Vanilla Conditional Flow Matching (Vanilla CFM) on a set of synthetic and experimental datasets. Our goal is to assess SP-FM’s generalizability and its robustness across multiple conditions, datasets, and hyperparameters. We seek empirical evidence of the need for the Gaussian Mixture Model (GMM) to describe the base distribution and its impact on the distributional distance between the inferred final distributions and the empirical ones.
5.1 SP-FM on Synthetic Data
@@
For SP-FM, we construct a mixture-conditioned base for each population. We set the number of base modes $I$ equal to the number of conditions and assign one Gaussian component per condition. In our synthetic setup, we do not use $h_{\bm{\Theta}}(\bm{y}_{i})$ to predict the component means; instead, the GMM mode locations are implemented as free learnable parameters, each initialized from a single rendered sample of its corresponding condition. The mixture weights are still predicted by the MLP $h_{\bm{p}}(\bm{y}_{i})$. The velocity field $v(\cdot,t;\bm{y}_{i})$ is then trained using the flow-matching objectives in Eqs. 67 to transport the resulting $\mu_{\mathrm{GMM}(\bm{\Theta}^{(i)},h_{\bm{p}}(\bm{y}_{i}))}$ toward $\rho_{i}$. As a baseline, conditional flow matching (CFM) fixes the base to $\mathcal{N}(\mathbf{0}_{D},\mathbf{I}_{D\times D})$ and samples $x_{0}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ at generation time.
@@
We train SP-FM and all baselines on a synthetic dataset of 6 letters with $R=20$ rotations per letter, and study rotation interpolation generalization. Specifically, we generate $R$ rotation angles uniformly spaced in $0,2\pi]$ and split them by index: the training set contains the 10 even-indexed rotations, while the held-out set contains the 10 odd-indexed rotations. We form the validation set by selecting a single val_letter and evaluating that letter only on the held-out (odd-indexed) rotations; all letters (including val_letter) are present in training but only at the even-indexed rotations. In the main paper we report results with $\texttt{val\_letter}=\texttt{S}$, and evaluate generalization on two unseen test letters, W and Y, each measured on their held-out (odd-indexed) rotations. To ensure conclusions are not tied to a particular character geometry, we repeat the same protocol with different choices of validation letter; results for $\texttt{val\_letter}=\texttt{H}$ are reported in the appendix. Results are reported in Table 1, with qualitative samples in Figure 2.
We compare SP-FM against three baselines: (i) CFM, which uses a fixed isotropic Gaussian base; (ii) a conditional GMM baseline that learns a mixture model by expectation-maximization (EM) algorithm with condition-dependent mixture weights $\pi(\cdot\mid\bm{y})$ (implemented as a learned lookup table over conditions) and shared component parameters, but does not learn a transport map; and (iii) SP-FM (K=1), an ablation where SP-FM is restricted to a single mixture component, reducing the base distribution to a unimodal Gaussian (learnable mean) and testing whether improved performance can be explained by a single adaptive mode alone. These baselines isolate the contributions of the two key ingredients in SP-FM: learning a condition-adaptive multi-modal base distribution and learning a conditional transport field.
For each setting, we select the best checkpoint based on the validation W2 distance and report all metrics at this checkpoint. Table 1 shows that SP-FM consistently outperforms all baselines across validation on S and testing on W and Y, demonstrating the effectiveness of mixture-conditioned flow matching for rotation-interpolation generalization. Importantly, the relative behavior of the baselines also highlights the contribution of each component: SP-FM (K=1) underperforms, indicating that a unimodal (single-mode) base is insufficient even when it is learnable, while the conditional GMM (EM) baseline underperforms SP-FM, indicating that adapting the base distribution alone (via $\pi(\cdot\mid\bm{y})$) is not enough without learning the subsequent transport field. Consistent with the training objective that explicitly encourages the predicted base to stay close to the target (Eq. 7), the Source vs. Target metrics are improved under SP-FM, providing direct empirical evidence that the learned source distribution is already similar to the target before transport. This behavior is also visible qualitatively in Figure 2: SP-FM starts from a source distribution that resembles the target shape and produces smoother, more faithful trajectories, whereas CFM exhibits larger mismatch at $t=0$ and more pronounced distortions along the path. Implementation details and hyperparameters for SP-FM and all baselines are provided in Appendix LABEL:sec:app:synthetic_details.
@@
Figure 2: Qualitative comparison of Conditional Flow Matching (CFM) and SP-FM on the synthetic letter–rotation benchmark. Each block shows samples at different integration times ($t=0$, $t=0.5$, $t=1$) and the corresponding target distribution. SP-FM produces smoother trajectories and more faithful reconstructions on unseen rotations, while CFM exhibits drift and shape distortion.
Table 1: Synthetic letter–rotation benchmark. Models are trained on even-indexed rotations and validated on S using held-out odd-indexed rotations; results are reported for S and unseen rotations on test letters W and Y. We report both Generated vs. Target and Source vs. Target distances. The conditional GMM baseline is trained via EM (no validation checkpoint), and since it has no flow model its Source vs. Target equals Generated vs. Target, so only the latter is reported.
Dataset Model Generated vs. Target Source vs. Target
MMD $\downarrow$ W1 $\downarrow$ W2 $\downarrow$ MMD $\downarrow$ W1 $\downarrow$ W2 $\downarrow$
Letter S (Val) CFM 0.01627 $\pm$ 0.00069 0.5874 $\pm$ 0.0199 0.7071 $\pm$ 0.0266 0.08249 $\pm$ 0.00138 1.0671 $\pm$ 0.0076 1.1898 $\pm$ 0.0069
SP-FM 0.00986 $\pm$ 0.00014 0.4740 $\pm$ 0.0028 0.5744 $\pm$ 0.0006 0.01986 $\pm$ 0.00008 0.6464 $\pm$ 0.0014 0.7388 $\pm$ 0.0005
SP-FM (I=1)
Letter W (Test) CFM 0.01809 $\pm$ 0.00081 0.6180 $\pm$ 0.0141 0.7283 $\pm$ 0.0152 0.06088 $\pm$ 0.00049 0.8972 $\pm$ 0.0062 1.0151 $\pm$ 0.0071
SP-FM (I=60) 0.01279 $\pm$ 0.00131 0.5323 $\pm$ 0.0203 0.6310 $\pm$ 0.0223 0.01825 $\pm$ 0.00062 0.6101 $\pm$ 0.0067 0.7126 $\pm$ 0.0061
GMM (I=600) 0.08931 $\pm$ 0.00387 1.1311 $\pm$ 0.0299 1.3202 $\pm$ 0.0375
SP-FM (I=1)
Letter Y (Test) CFM 0.02333 $\pm$ 0.00197 0.6278 $\pm$ 0.0160 0.7415 $\pm$ 0.0176 0.06020 $\pm$ 0.00128 1.0614 $\pm$ 0.0048 1.2410 $\pm$ 0.0054
SP-FM (I=60) 0.01799 $\pm$ 0.00015 0.5433 $\pm$ 0.0044 0.6403 $\pm$ 0.0064 0.04079 $\pm$ 0.00225 0.9336 $\pm$ 0.0297 1.0861 $\pm$ 0.0312
GMM (I=600) 0.07423 $\pm$ 0.00528 0.8986 $\pm$ 0.0317 1.0508 $\pm$ 0.0417
SP-FM (I=1)
5.2 Using SP-FM to predict single-cell responses for unseen interventions